OpenStudy (anonymous):

A company manufactures tote bags. The company spends $1200 each day for overhead expenses plus $9 per tote bag for labor and materials. The toe bags sell for $25 each. How many tote bags must the company sell each day to equal its daily costs for overhead, labor and materials(break even point)? Write a equation and solve.

OpenStudy (anonymous):

1200+9X=25X OR 1200=25X-9X OR 1200=16X OR X=1200/16 OR X=75 BAGS TO BREAK EVEN. PROOF 1200+9*75=25*75 OR 1200+675=1875 OR 1875=1875 MEDAL?????? PLEASE???

OpenStudy (anonymous):

yes!!!!